# Weighted Fisher divergence for high-dimensional Gaussian variational inference

### v1

@@ 2 Ordering of divergences for
We first prove $\Sigma^{\text{M}}_{ii}\leq\Sigma^{\text{KL}}_{ii}$ $\forall\,i$. From (1),
$\Sigma^{\text{M}}_{ii}=\sqrt{\frac{M_{ii}}{\sum_{j=1}^{d}M_{jj}\Lambda_{ij}^{2}}}\leq\sqrt{\frac{M_{ii}}{M_{ii}\Lambda_{ii}^{2}}}=\frac{1}{\Lambda_{ii}}=\Sigma_{ii}^{\text{\rm KL}}.$ (5)
Since $\Lambda$ has at least one nonzero off-diagonal entry, $\exists\,i\in\{1,\dots,d\}$ such that the inequality in (5) is strict. The proof for $\Sigma^{\text{S}}_{ii}\leq\Sigma^{\text{KL}}_{ii}$ is given in Margossian et al. (2024) and we include it here for entirety. From the KKT conditions discussed earlier, if case (a) applies, then $\Sigma^{\text{S}}_{ii}=0<\Sigma^{\text{KL}}_{ii}$. Otherwise, case (b) applies and (4) implies that
$\Lambda_{ii}^{2}\Sigma^{\text{S}}_{ii}\leq\sum_{j}\Lambda_{ij}^{2}\Sigma^{\text{S}}_{ii}=\Lambda_{ii}\implies\Sigma^{\text{S}}_{ii}\leq\frac{1}{\Lambda_{ii}}=\Sigma^{\text{KL}}_{ii}.$ (6)
To obtain the strict inequality, note that if case (a) applies for at least one $i$, then $\Sigma_{ii}^{S}<\Sigma_{ii}^{KL}$ for such an $i$. Otherwise, case (b) applies $\forall$ $i$. Since $\Lambda$ has at least one nonzero off-diagonal entry, $\exists\,i\in\{1,\dots,d\}$ such that the first inequality in (6) is strict. ∎
From Theorem 1, both the weighted Fisher and score-based divergences tend to underestimate the posterior variance more severely than KLD under the mean-field assumption, but ordering between the Fisher and score-based divergences is more nuanced. If case (a) of the KKT conditions apply, then $\Sigma^{\text{S}}_{ii}=0<\Sigma^{\text{F}}_{ii}$. If case (b) applies, then from (2) and (4),
$\Lambda_{ii}^{2}\Sigma^{\text{S}}_{ii}$ $\leq\sum_{j=1}^{d}\Lambda_{ij}^{2}\Sigma^{\text{S}}_{jj}=\Lambda_{ii}\Sigma^{\text{F}}_{ii}\sqrt{\sum\nolimits_{j=1}^{d}\Lambda_{ij}^{2}}$
$\implies\Sigma^{\text{S}}_{ii}\leq\Sigma^{\text{F}}_{ii}\frac{\sqrt{\sum_{j=1}^{d}\Lambda_{ij}^{2}}}{\Lambda_{ii}}.$
@@
$\Sigma^{\text{S}}_{ii}$ $\leq\Sigma^{\text{F}}_{ii}\frac{\sqrt{\Lambda_{ii}^{2}+\sum_{j\neq i}\Lambda_{ij}^{2}}}{\Lambda_{ii}}\leq\Sigma^{\text{F}}_{ii}\frac{\sqrt{\Lambda_{ii}^{2}+(\sum_{j\neq i}|\Lambda_{ij}|)^{2}}}{\Lambda_{ii}}$
$\leq\Sigma^{\text{F}}_{ii}\frac{\sqrt{\Lambda_{ii}^{2}+\Lambda_{ii}^{2}}}{\Lambda_{ii}}=\sqrt{2}\Sigma^{\text{F}}_{ii}.$
Thus the ratio of $\Sigma^{\text{S}}_{ii}/\Sigma^{\text{F}}_{ii}$ is bounded by $\sqrt{2}$ $\forall\;i$ if $\Lambda$ is diagonally dominant.
For a more concrete comparison of posterior variance approximation based on Fisher and score-based divergences, consider a three-dimensional Gaussian target with precision matrix,
@@
For Fisher divergence, $\Sigma_{ii}^{\text{F}}$ can be obtained from (2), while the splitting conic solver (SCS, O’Donoghue et al., 2016) in the CVXPY Python package is used to solve the NQP in (3) for score-based divergence. SCS is designed for convex optimization problems characterized by conic constraints, such as nonnegative constraints. It decomposes the optimization task into smaller subproblems solved iteratively by operator-splitting techniques.
Fig 1 illustrates how variance parameters obtained from the Fisher and score-based divergences compare by varying the conditional correlations $a$, $b$ and $c$, with each plot representing a value of $c$. The colored regions represent configurations for which $\Lambda$ is positive definite. The red, blue and green regions indicate where $\Sigma_{ii}^{S}\leq\Sigma_{ii}^{F}$ for all cases, only two cases and only one case respectively. There is no region such that $\Sigma_{ii}^{S}>\Sigma_{ii}^{F}$ $\forall$ $i$. Variance estimates based on score-based divergence have a higher tendency to exceed those based on Fisher divergence when the conditional correlation $a$, $b$ or $c$ has a large magnitude. The orange-bordered region indicates where $\Lambda$ is diagonally dominant, and $\Sigma^{\text{S}}_{ii}/\Sigma^{\text{F}}_{ii}$ can be bounded even more tightly by 1 instead of only $\sqrt{2}$ in this example.
@@ 3 Ordering of divergences for non-Gaussian target
Next, we study ordering of the Fisher, score-based and KL divergences in posterior mode and variance estimation when the target $p(\theta|y)$ is a univariate non-Gaussian density, while the variational density $q(\theta)$ is ${\text{N}}(\mu,\sigma^{2})$ with mean $\mu\in\mathbb{R}$ and standard deviation $\sigma>0$. In the examples below, the VI objective function is available in closed form or can be computed numerically, and optimal variational approximations are obtained using L-BFGS via the optim function in R. For univariate densities, $S(q\|p)=\sigma^{2}F(q\|p)$, so we only present the Fisher divergence. All derivations are given in the supplement.
Let $m_{*}$ and $\sigma^{2}_{*}$ denote the mode and variance of the target density. To evaluate the performance of different divergences, we use three criteria, normalized absolute difference in mode: $|\mu-m_{*}|/\sigma_{*}$, variance ratio: $\sigma^{2}/\sigma^{2}_{*}$, and integrated absolute error : $\text{IAE}(q)=\int|q(\theta)-p(\theta|y)|d\theta$, which is invariant under monotone transformations of $\theta$ and lies in [0, 2]. We define $\text{Accuracy}(q)=1-\text{IAE}(q)/2$, where a higher accuracy indicates a better approximation of the target.
@@ 3.1 Student’s $t$
First, we consider the univariate Student’s $t$ as target, which is symmetric but has heavier tails than a Gaussian. The probability density function (pdf) for $\theta\sim t(\nu)$ is $p(\theta|y)=\left(1+\theta^{2}/\nu\right)^{-(\nu+1)/2}\Gamma(\frac{\nu+1}{2})/(\sqrt{\pi\nu}\,\Gamma(\frac{\nu}{2}))$, where $\nu\in\{3,5,10\}$ is the degrees of freedom. For KLD, we maximize the lower bound,
${\mathcal{L}}$ $=\log\left\{\Gamma\left(\frac{\nu+1}{2}\right)\right\}-\frac{1}{2}\log\left\{\Gamma^{2}\left(\frac{\nu}{2}\right)\nu\right\}$
The Fisher divergence is
$F(q\|p)$ $=(\nu+1)^{2}{\text{E}}_{q}\left\{\frac{\theta^{2}}{(\nu+\theta^{2})^{2}}\right\}+\frac{1}{\sigma^{2}}$
All three divergences successfully capture the mode of the target at 0. From Table 1, SD exhibits the most posterior variance underestimation, followed by FD, while KLD has the least underestimation. In terms of the IAE, both FD and SD yield approximations with higher accuracy than KLD. Fig 2 (first 3 plots) compares variational densities with the target. KLD tends to underestimate the probability mass around the mode more severely than FD and SD.
@@
Table 1: Variance ratio and accuracy for Student’s $t$.
Figure 3: Gaussian variational approximations for skew normal.
$|\mu-m_{*}|/\sigma_{*}$ $\sigma^{2}/\sigma^{2}_{*}$ accuracy
$(t,\lambda)$ KLD FD SD KLD FD SD KLD FD SD
(1, 1) 0.07 0.07 0.07 0.992 0.984 0.979 98.27 98.31 98.32
(1, 2) 0.25 0.23 0.20 0.919 0.851 0.803 93.77 93.81 93.79
(1, 5) 0.66 0.91 0.07 0.677 0.642 0.248 83.93 76.44 68.50
(5, 1) 0.66 0.91 0.07 0.677 0.642 0.248 83.92 76.42 68.49
(5, 2) 0.94 2.20 0.10 0.504 0.757 0.054 76.50 45.38 37.06
(5, 5) 1.20 2.94 0.09 0.352 0.644 0.008 68.00 30.35 16.35
Table 2: Results of KL, Fisher and score-based divergences for skew normal.
3.2 Log transformed inverse-gamma
Next, consider the normal sample model in Tan and Chen (2024), where $y_{i}|\theta\sim{\text{N}}(0,\exp(\theta))$ for $i=1,\dots,n$, with prior $\exp(\theta)\sim{\text{IG}}(a_{0},b_{0})$ and $a_{0}=b_{0}=0.01$. The true posterior of $\exp(\theta)$ is ${\text{IG}}(a_{1},b_{1})$, where $a_{1}=a_{0}+n/2$ and $b_{1}=b_{0}+\sum_{i=1}^{n}y_{i}^{2}/2$. We simulate $n=6$ observations by setting $\exp(\theta)=225$. The lower bound for minimizing the KLD is
${\mathcal{L}}$ $=\frac{1-n}{2}\log(2\pi)+a_{0}\log b_{0}-\log\Gamma(a_{0})-a_{1}\mu$
$\quad-\exp\left(\frac{\sigma^{2}-2\mu}{2}\right)T+\frac{1}{2}\log(\sigma^{2})+\frac{1}{2}.$
The Fisher divergence is
$F(q\|p)$ $=a_{1}^{2}+b_{1}^{2}\exp\left(2\sigma^{2}-2\mu\right)$
$\quad-2b_{1}\left(a_{1}+1\right)\exp(\sigma^{2}/2-\mu)+{1}/{\sigma^{2}}.$
Table 3 shows that SD provides the most accurate mode approximation but also exhibits the most severe variance underestimation. On the other hand, KLD yields the least accurate mode approximation while most closely matching the target variance. Overall, KLD has the highest accuracy followed by FD and SD. A visualization is given in Fig 2 (last plot).
@@ 3.2 Log transformed inverse-gamma
Table 3: Results for log transformed inverse-gamma.
@@ 3.3 Skew normal
Finally, let the target be a univariate skew normal. The pdf of $\theta\sim{\text{SN}}(m,t,\lambda)$ is $p(\theta|y)=2\,\phi(\theta|m,t^{2})\,\Phi\{\lambda(\theta-m)\}$, where $m\in\mathbb{R}$, $t>0$ and $\lambda\in\mathbb{R}$ are the location, scale and skewness parameters respectively, and $\Phi(\cdot)$ is cumulative distribution function (cdf) of the standard normal. For KLD, we maximize the lower bound,
${\mathcal{L}}$ $=\log 2-\log(t)-\frac{\sigma^{2}+(\mu-m)^{2}}{2t^{2}}$
The Fisher divergence is
$\quad-\frac{2\lambda}{t^{2}}\,{\text{E}}_{q}\left[(\theta-m)\frac{\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}\right]-\frac{2}{t^{2}}$
$\quad+\frac{2\lambda}{\sigma^{2}}{\text{E}}_{q}\left[(\theta-\mu)\frac{\phi\{\lambda(\theta-m)\}}{\Phi\{\lambda(\theta-m)\}}\right]+\frac{1}{\sigma^{2}}.$
We set $m=0$ and let $t\in\{1,5\}$ and $\lambda\in\{1,2,5\}$. From Table 2, SD captured the mode most accurately across all settings, followed by KLD and FD (for which estimation of the mode is very poor when both scale and skewness are large). All three divergences underestimated the variance, with KLD and FD exhibiting less underestimation than SD (for which the variance estimate approaches zero as $t$ and $\lambda$ increase). The accuracies of FD and SD are very close to KLD when $t=1$ and $\lambda\in\{1,2\}$. However, KLD has higher accuracy than both FD and SD as skewness and scale increase. Fig 3 shows density plots of the target and optimal variational approximations. SD is good at identifying the mode, whereas FD and KLD estimate the variance more accurately. Multiple local minimums were detected for the SD in this context.
The examples in this section suggest that SD captures the mode more accurately than FD and KLD if the target density is skewed, but underestimates the posterior variance most severely. KLD has higher accuracy than both FD and SD when the target is skewed, but lower accuracy when the target is symmetric and has heavy tails.

### v2

@@ 2 Ordering of divergences for Gaussian target
We first prove $\hat{\Sigma}^{\text{M}}_{ii}\leq\hat{\Sigma}^{\text{KL}}_{ii}$ $\forall\,i$. From (1),
$\hat{\Sigma}^{\text{M}}_{ii}=\sqrt{\frac{M_{ii}}{\sum_{j=1}^{d}M_{jj}\Lambda_{ij}^{2}}}\leq\sqrt{\frac{M_{ii}}{M_{ii}\Lambda_{ii}^{2}}}=\frac{1}{\Lambda_{ii}}=\hat{\Sigma}_{ii}^{\text{\rm KL}}.$ (5)
Since $\Lambda$ has at least one nonzero off-diagonal entry, $\exists\,i\in\{1,\dots,d\}$ such that the inequality in (5) is strict. The proof for $\hat{\Sigma}^{\text{S}}_{ii}\leq\hat{\Sigma}^{\text{KL}}_{ii}$ is given in Margossian et al. (2024) and we include it here for entirety. From the KKT conditions discussed earlier, if case (a) applies, then $\hat{\Sigma}^{\text{S}}_{ii}=0<\hat{\Sigma}^{\text{KL}}_{ii}$. Otherwise, case (b) applies and (4) implies that
$\Lambda_{ii}^{2}\hat{\Sigma}^{\text{S}}_{ii}\leq\sum_{j}\Lambda_{ij}^{2}\hat{\Sigma}^{\text{S}}_{jj}=\Lambda_{ii}\implies\hat{\Sigma}^{\text{S}}_{ii}\leq\frac{1}{\Lambda_{ii}}=\hat{\Sigma}^{\text{KL}}_{ii}.$ (6)
To obtain the strict inequality, note that if case (a) applies for at least one $i$, then $\hat{\Sigma}_{ii}^{S}<\hat{\Sigma}_{ii}^{KL}$ for such an $i$. Otherwise, case (b) applies $\forall$ ii. Since $\Lambda$ has at least one nonzero off-diagonal entry, $\exists\,i\in\{1,\dots,d\}$ such that the first inequality in (6) is strict. ∎
From Theorem 1, both the weighted Fisher and score-based divergences tend to underestimate the posterior variance more severely than KLD under mean-field, but the ordering between FD and SD is more nuanced. If case (a) of the KKT conditions apply, then $\hat{\Sigma}^{\text{S}}_{ii}=0<\hat{\Sigma}^{\text{F}}_{ii}$. If case (b) applies, then from (2) and (4),
$\Lambda_{ii}^{2}\hat{\Sigma}^{\text{S}}_{ii}$ $\leq\sum_{j=1}^{d}\Lambda_{ij}^{2}\hat{\Sigma}^{\text{S}}_{jj}=\Lambda_{ii}\hat{\Sigma}^{\text{F}}_{ii}\sqrt{\sum\nolimits_{j=1}^{d}\Lambda_{ij}^{2}}$
$\implies\hat{\Sigma}^{\text{S}}_{ii}\leq\hat{\Sigma}^{\text{F}}_{ii}\frac{\sqrt{\sum_{j=1}^{d}\Lambda_{ij}^{2}}}{\Lambda_{ii}}.$
@@
$\hat{\Sigma}^{\text{S}}_{ii}$ $\leq\hat{\Sigma}^{\text{F}}_{ii}\frac{\sqrt{\Lambda_{ii}^{2}+\sum_{j\neq i}\Lambda_{ij}^{2}}}{\Lambda_{ii}}\leq\hat{\Sigma}^{\text{F}}_{ii}\frac{\sqrt{\Lambda_{ii}^{2}+(\sum_{j\neq i}|\Lambda_{ij}|)^{2}}}{\Lambda_{ii}}$
$\leq\hat{\Sigma}^{\text{F}}_{ii}\frac{\sqrt{\Lambda_{ii}^{2}+\Lambda_{ii}^{2}}}{\Lambda_{ii}}=\sqrt{2}\hat{\Sigma}^{\text{F}}_{ii}.$
Thus the ratio of $\hat{\Sigma}^{\text{S}}_{ii}/\hat{\Sigma}^{\text{F}}_{ii}$ is bounded by $\sqrt{2}$ $\forall\;i$ if $\Lambda$ is diagonally dominant.
For a more concrete comparison of posterior variance approximation based on FD and SD, consider a three-dimensional Gaussian target with precision matrix,
@@
For FD, $\hat{\Sigma}_{ii}^{\text{F}}$ can be obtained from (2), while the splitting conic solver (SCS, O’Donoghue et al., 2016) in the CVXPY Python package is used to solve the NQP in (3) for SD. SCS is designed for convex optimization problems characterized by conic constraints, such as non-negativity. It decomposes the optimization into subproblems solved iteratively by operator-splitting techniques.
Fig 1 illustrates how variance parameters obtained from FD and SD compare by varying the conditional correlations $a$, $b$ and $c$. Each plot represents a value of $c$. The colored regions represent configurations for which $\Lambda$ is positive definite, and there is no region where $\hat{\Sigma}_{ii}^{S}>\hat{\Sigma}_{ii}^{F}$ $\forall$ ii. Variance estimates based on SD are more likely to exceed those based on FD when $a$, $b$ or $c$ has a large magnitude. In this example, $\hat{\Sigma}^{\text{S}}_{ii}/\hat{\Sigma}^{\text{F}}_{ii}$ can be bounded more tightly by 1 instead of only $\sqrt{2}$ when $\Lambda$ is diagonally dominant.
@@ 3 Ordering of divergences for non-Gaussian target
Next, we study the ordering of FD, SD and KLD in posterior mean, mode and variance estimation when the target distribution is non-Gaussian while the variational approximation is Gaussian. Theoretical analysis in this setting is complex and numerical methods are often required. We show that the true posterior mean is recoverable across all divergences for the multivariate Student’s $t$, and an ordering of the mean, mode and variance estimation is established for the log transformed inverse gamma density. Otherwise, empirical comparisons are made by considering $p(\theta\mid y)$ as some univariate non-Gaussian density, and the variational density $q(\theta)$ as ${\text{N}}(\mu,\sigma^{2})$. Results in this section indicate that KLD estimates the mean most accurately and has the highest accuracy (as defined below) when the target is skewed, but the lowest accuracy when the target is symmetric and has heavy tails. SD captures the mode most accurately if the target density is skewed, but underestimates the posterior variance most severely.
Let $\mu_{*}$, $m_{*}$ and $\sigma^{2}_{*}$ denote the mean, mode and variance of the target density. To evaluate the performance of different divergences, we use the normalized absolute difference in mean and mode: $|\mu-\mu_{*}|/\sigma_{*}$ and $|\mu-m_{*}|/\sigma_{*}$, variance ratio: $\sigma^{2}/\sigma^{2}_{*}$, and integrated absolute error : $\text{IAE}(q)=\int|q(\theta)-p(\theta|y)|d\theta\in[0,2]$, which is invariant under monotone transformations of $\theta$. We define $\text{accuracy}(q)=1-\text{IAE}(q)/2$, where a higher value indicates a more accurate approximation of the target. In the examples below, the VI objective function is tractable or can be computed numerically, and variational approximations are optimized using L-BFGS via optim in R.
@@ 3.1 Student’s $t$
First, consider the multivariate Student’s $t$ distribution, $t_{\nu}(m,S)$ for $\theta\in\mathbb{R}^{d}$ as the target, where
$p(\theta\mid y)=\frac{\Gamma\left(\frac{\nu+d}{2}\right)\left(1+(\theta-m)^{\top}S^{-1}(\theta-m)/\nu\right)^{-\frac{\nu+d}{2}}}{\Gamma\left(\frac{\nu}{2}\right)\,(\nu\pi)^{d/2}\,|S|^{1/2}}.$
The Student’s $t$ is symmetric but has heavier tails than the Gaussian, and $\nu$, $m$ and $S$ denote the degrees of freedom, location parameter and scale matrix respectively. Theorem 2 shows that the true posterior mean or mode of the Student’s $t$ is recoverable by a Gaussian variational approximation under all three divergences.
Theorem 2.
Let $q(\theta)={\text{N}}(\mu,\Sigma)$ and $p(\theta\mid y)=t_{\nu}(m,S)$. Then $\mu=m$ is a stationary point of the KLD, FD and SD between $q(\theta)$ and $p(\theta\mid y)$.
Next, we consider the univariate Student’s $t$ as the target to compare the performances of different divergences in capturing the variance. For $\theta\sim t(\nu)$, $p(\theta|y)=\left(1+\theta^{2}/\nu\right)^{-(\nu+1)/2}\Gamma(\frac{\nu+1}{2})/(\sqrt{\pi\nu}\,\Gamma(\frac{\nu}{2}))$, where $\nu\in\{3,5,10\}$ is the degrees of freedom. All divergences successfully capture mode of the target at 0, verifying Theorem 2. From Table 1, SD exhibits the most severe posterior variance underestimation, followed by FD and then KLD. In terms of the IAE, both FD and SD yield approximations with higher accuracy than KLD. Fig 2 (first 3 plots) compares optimal variational densities with the target, and showing that KLD tends to underestimate the mass around the mode more severely than FD and SD.
@@
Table 1: Results for Student’s $t$ (best values highlighted in bold).
3.2 Log transformed inverse gamma
Consider the normal sample model in Tan and Chen (2024), where $y_{i}\mid\theta\sim{\text{N}}(0,\exp(\theta))$ for $i=1,\dots,n$, with prior, $\exp(\theta)\sim{\text{IG}}(a_{0},b_{0})$, and $a_{0}=b_{0}=0.01$. The true posterior of $\exp(-\theta)$ is $\text{G}(a_{1},b_{1})$, where $a_{1}=a_{0}+n/2$ and $b_{1}=b_{0}+\sum_{i=1}^{n}y_{i}^{2}/2$. The true posterior mode, mean and variance of $\theta$ are $m_{*}=\log(b_{1}/a_{1})$, $\mu_{*}=\log b_{1}-\psi(a_{1})$ and $\sigma_{*}^{2}=\psi_{1}(a_{1})$, where $\psi(\cdot)$ and $\psi_{1}(\cdot)$ denote the digamma and trigamma functions respectively.
This is a rare example where the FD, SD and evidence lower bound for the KLD can be derived in closed form. Moreover, the optimal variational parameters for all three divergences are available analytically, as given in Table 2. Note that $W_{0}(\cdot)$ denotes the principal branch of the Lambert W function (Corless et al., 1996). Theorem 3 shows that SD underestimates the variance most severely, followed by FD and then KLD. Moreover, SD yields the best estimate of the mode, while KLD estimates the mean most accurately, with FD lying in between.
KLD $\hat{\sigma}_{\text{\rm KL}}^{2}=\frac{1}{a_{1}}$ $\hat{\mu}_{\text{\rm KL}}=\log\frac{b_{1}}{a_{1}}+\frac{1}{2a_{1}}$
FD $\hat{\sigma}^{2}_{\text{F}}=-2W_{0}\left(-\frac{1}{2(a_{1}+1)}\right)$ $\hat{\mu}_{\text{F}}=\log\frac{b_{1}}{a_{1}+1}+\frac{3\hat{\sigma}^{2}_{\text{F}}}{2}$
SD $\hat{\sigma}^{2}_{\text{S}}=1-W_{0}\left(\frac{{\rm e}a_{1}^{2}}{(a_{1}+1)^{2}}\right)$ $\hat{\mu}_{\text{S}}=\log\frac{b_{1}}{a_{1}+1}+\frac{3\hat{\sigma}^{2}_{\text{S}}}{2}$
Table 2: Optimal variational parameters for log transformed inverse gamma.
Theorem 3.
Let $\hat{\mu}_{\text{\rm KL}}$, $\hat{\mu}_{\text{F}}$, $\hat{\mu}_{\text{S}}$, $\hat{\sigma}^{2}_{\text{\rm KL}}$, $\hat{\sigma}^{2}_{\text{F}}$ and $\hat{\sigma}^{2}_{\text{S}}$ denote the optimal mean and variance parameters that minimize the KLD, FD and SD respectively, when the target is a log transformed inverse gamma density and the variational approximation is Gaussian. Then
$\hat{\sigma}^{2}_{\text{S}}<\hat{\sigma}^{2}_{\text{F}}<\hat{\sigma}^{2}_{\text{\rm KL}}<\sigma_{*}^{2},$
$m_{*}<\hat{\mu}_{\text{S}}<\hat{\mu}_{\text{F}}<\hat{\mu}_{\text{\rm KL}}<\mu_{*},$
where $\mu_{*}$, $m_{*}$ and $\sigma^{2}_{*}$ denote the mean, mode and variance of the target.
To verify Theorem 3, we simulate $n=6$ observations by setting $\exp(\theta)=225$. Table 3 shows that the ordering in mean, mode and variance estimation is consistent with Theorem 3. Overall, KLD has the highest accuracy followed by FD and then SD. A visualization is given in Fig 2 (last plot).
@@ 3.2 Log transformed inverse gamma
Table 3: Results for log transformed inverse gamma (best values highlighted in bold).
$|\mu-\mu_{*}|/\sigma_{*}$ $|\mu-m_{*}|/\sigma_{*}$ $\sigma^{2}/\sigma^{2}_{*}$ accuracy
$(t,\lambda)$ KLD FD SD KLD FD SD KLD FD SD KLD FD SD
(1, 1) 0.001 0.003 0.004 0.070 0.067 0.066 0.992 0.984 0.979 98.27 98.31 98.32
(1, 2) 0.006 0.031 0.064 0.255 0.230 0.197 0.919 0.851 0.803 93.77 93.81 93.79
(1, 5) 0.004 0.251 0.586 0.657 0.912 0.075 0.677 0.642 0.248 83.93 76.44 68.50
(5, 1) 0.004 0.251 0.586 0.657 0.912 0.075 0.677 0.642 0.248 83.92 76.42 68.49
(5, 2) 0.024 1.285 1.011 0.939 2.200 0.097 0.504 0.757 0.054 76.50 45.38 37.06
(5, 5) 0.077 1.819 1.209 1.201 2.942 0.086 0.352 0.644 0.008 68.00 30.35 16.35
Table 4: Results for skew normal (best values highlighted in bold).
Figure 3: Gaussian variational approximations for skew normal.
@@ 3.3 Skew normal
Finally, let the target be a univariate skew normal, $\theta\sim{\text{SN}}(m,t,\lambda)$. Then $p(\theta|y)=2\,\phi(\theta|m,t^{2})\,\Phi\{\lambda(\theta-m)\}$, where $m\in\mathbb{R}$, $t>0$ and $\lambda\in\mathbb{R}$ are the location, scale and skewness parameters respectively, and $\Phi(\cdot)$ is cumulative distribution function of the standard normal.
We set $m=0$ and let $t\in\{1,5\}$ and $\lambda\in\{1,2,5\}$. From Table 4, KLD estimates the mean most accurately, while SD captures the mode most accurately. For FD, estimation of the mode is very poor when both scale and skewness are large. SD underestimates the variance most severely, with the variance estimate collapsing to zero as $t$ and $\lambda$ increase. KLD has higher accuracy than both FD and SD as skewness and scale increase. From Fig 3, SD is good at identifying the mode, whereas FD and KLD estimate the variance more accurately. We note that multiple local minimums were detected for SD in this context.